Given two numbers that have a
difference of 7 and add to make -1

Answer by Edwin McCravy(20056)

>>Give two numbers...<<
Let them be x and y
>>that have a difference of 7...<<
x-y=7
>>...and add to make -1<<
x+y=-1

Add the two equations term by term

 x-y= 7
 x+y=-1
-------
2x  = 6
 x  = 3

Substitute 3 for x in x-y=7

x-y=7
3-y=7
 -y=4
  y=-4

They are 3 and -4.

They have a difference of 7 because 3-(-4) = 3+4 = 7

They add to make -1 because 3+(-4) = -1

Edwin